# Retrieval-Augmented Generation with Estimation of Source Reliability

### v2

@@ 5 Experiments
Figure 2: Accuracy performance on NQ and TQA datasets. (a) Results with heterogeneous reliability via beta priors for varying sources (4 to 9) across the Llama3-8B-Instruct and GPT-4o-mini models. See Appendix E.1 on the HotpotQA dataset and Phi3-mini-Instruct model. (b) Results with adversarial setting via adversary-hammer prior for varying adversaries (1 to 7) with Llama3-8B-Instruct model, highlighting overall trends. Exact values, which may overlap significantly, are provided in Appendix E.2 with HotpotQA results.
@@ 5.1 Experimental setups
Datasets. We create a multi-source RAG dataset with heterogeneous source reliability, using three question-answering (QA) datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQA) (Joshi et al., 2017), and HotpotQA (Yang et al., 2018). For each dataset, we generate both diverse factual documents and misinformation. Each source $\mathcal{S}_{i}$ is characterized by two parameters: reliability $p_{i}$, which represents the probability of providing factual information, and coverage $r_{i}$, which indicates the probability of containing relevant documents for a given query. To model source reliability $p_{i}$, we adopt two widely used priors from the reliability estimation literature (Liu et al., 2012; Li & Yu, 2014):
  • Beta prior: $p_{i}$ is sampled from $\text{Beta}\left(\nicefrac{{2\bar{w}}}{{1-\bar{w}}},\ 2\right)$ with an expected mean of $\bar{w}$. This setup reflects scenarios where sources exhibit a continuous spectrum of reliability, rather than strictly “reliable” or “unreliable”. Following Liu et al. (2012); Li & Yu (2014), we set $\bar{w}=0.6$, balancing the presence of reliable and unreliable sources.
@@
For analytical simplicity, we set $r_{i}=0.6$ for both priors to focus on evaluating $p_{i}$. The details of the data generation and source construction processes are provided in Appendix G. Due to the computational and financial constraints, we use 1,600 queries per dataset, allocating 200 queries for reliability estimation and 1,400 queries for test evaluation.
Baselines. We compare our framework against RA-RAG and six baselines. (1) Oracle WMV assumes perfect knowledge of source reliability and directly uses these values as weights in Equation (1), representing the ideal scenario for multi-source RAG. (2) MV assigns equal weight to all sources, setting $v_{i}=1$ in Equation (1), disregarding source reliability. (3) Vanilla RAG (Lewis et al., 2020) follows the standard RAG approach, retrieving documents without additional modules. (4) Robust RAG (Xiang et al., 2024) is the first certifiably robust defense framework that enhances robustness by aggregating keywords from independent passages, assuming that the majority of retrieved documents are trustworthy. (5) Self-RAG (Asai et al., 2024) is an advanced RAG that improves performance through adaptive retrieval, reducing irrelevant documents by leveraging specialized reflection tokens to improve factual accuracy. (6) Vanilla LLM generates responses without retrieval. Among these baselines, (1) and (2) are designed for multi-source RAG, while (3), (4), and (5) follow the standard RAG approach.
Models. For language models, we use Llama3-8B-Instruct (Dubey et al., 2024), Phi3-mini-Instruct (Abdin et al., 2024), GPT-4o-mini (OpenAI, 2024), and Llama2-7B (Touvron et al., 2023). As a retriever, we use Contriever (Izacard et al., 2022). Due to space limitations, the results for Llama2-7B with Self-RAG fine-tuned on Llama2-7B are provided in Appendix D.
@@
Figure 3: A qualitative example comparing the answers produces by MV and RA-RAG for a query from the NQ dataset. Additional examples are available in Appendix F.
Evaluation metric. Following prior works (Mallen et al., 2023; Asai et al., 2024), we use accuracy as an evaluation metric, based on whether gold answers are included in model-generated responses. All results are averaged over $10$ random trials.
@@ 5.2 Main results
Beta prior. We evaluate RA-RAG across varying numbers of sources to assess its effectiveness in heterogeneous source reliability. As shown in Figure 2(a), RA-RAG consistently outperforms baselines, with performance gains increasing as more sources are incorporated. These results demonstrate the robustness of our approach in aggregating information from multiple sources with varying reliability. Notably, by selecting a subset of reliable and relevant sources using $\kappa$-RRSS, RA-RAG achieves performance comparable to Oracle WMV while improving efficiency by relying on fewer sources. In contrast, Robust RAG struggles with varying source reliability, as its certification assumption does not hold, resulting in lower performance than MV. Additionally, RA-RAG significantly outperforms Self-RAG, as shown in the Appendix D. These results emphasize the importance of differentiating between sources to prevent retrieval results from being overwhelmed by misinformation.
Figure 3 highlights the importance of considering source reliability when aggregating information across sources. While MV selects “November 2018” based only on response frequency, although it has low reliability, RA-RAG correctly identifies “November 2017” by leveraging well-estimated source reliabilities.
Adversary-hammer prior. To evaluate the robustness of RA-RAG in the worst-case scenario, we use the adversary-hammer prior with a total of 9 sources on the NQ dataset with Llama3-8B-Instruct, as shown in Figure 2(b). Our RA-RAG demonstrates significant robustness against adversaries, whereas Robust RAG and Vanilla RAG suffer severe performance degradation as the number of adversaries increases. Similarly, Self-RAG experiences significant performance degradation, as detailed in Appendix D. Notably, when the number of adversaries exceeds four, the performance of MV significantly degrades due to the dominance of misinformation, leading MV to select incorrect answers.
Table 1: Comparison of computational efficiency with and without $\kappa$-RRSS across various computational metrics (token consumption, API calls, and inference cost) and accuracy, evaluated on the NQ dataset using GPT-4o-mini. The reported values represent the average per query, with the values in parentheses ()($\cdot$)( ⋅ ) indicating the reduction rate achieved with $\kappa$-RRSS.
# Src $\kappa$-RRSS Token Consumption ($\downarrow$) API Calls ($\downarrow$) Inference Cost ($\downarrow$) Accuracy ($\uparrow$) $5$ w/o $3138$ $5$ $0.00048$ $0.597$ w/ $3055$ ($\downarrow 2.6\%$) $4.87$ ($\downarrow 2.6\%$) $0.00046$ ($\downarrow 4.2\%$) $0.597$ $10$ w/o $6272$ $10$ $0.00096$ $0.744$ w/ $4251$ ($\downarrow 32.3\%$) $6.79$ ($\downarrow 32.1\%$) $0.00065$ ($\downarrow 32.3\%$) $0.727$ $20$ w/o $12578$ $20$ $0.00192$ $0.769$ w/ $4644$ ($\downarrow 63.1\%$) $7.42$ ($\downarrow 62.9\%$) $0.00071$ ($\downarrow 63.0\%$) $0.748$
Figure 4: Accuracy across different values of $\kappa$ on Llama3-8B-Instruct and NQ dataset. Results for other datasets are provided in Appendix E.4
@@ 5.3 Ablation studies and analysis
Impact of $\kappa$ for $\kappa$-RRSS. We conduct an ablation study on $\kappa$ across three datasets using Llama3-8B-Instruct with 9 sources. As shown in Figure 4, RA-RAG achieves stable performance starting from $\kappa=4$, indicating that selecting a small subset of reliable and relevant sources can maintain performance while significantly reducing computational overhead. This trend is consistent across other datasets; refer to Appendix E.4.
Computational efficiency of $\kappa$-RRSS. To investigate the impact of $\kappa$-RRSS on computational efficiency, we compare RA-RAG in two configurations: with and without $\kappa$-RRSS. We measure three computational metrics: token consumption, API calls, and inference cost. Specifically, token consumption refers to the total number of tokens processed per query during inference, including both input and output tokens. API calls measure the number of external API requests per query. Inference cost represents the computational expense (\$ per query) based on the GPT-4o-mini pricing policy.
As shown in Table 1, incorporating $\kappa$-RRSS consistently enhances computational efficiency across all metrics, with the reduction rate increasing as the number of sources grows. For example, in terms of token consumption, $\kappa$-RRSS reduces the total tokens processed by 2.6% with 5 sources, 32.3% with 10 sources, and 63.1% with 20 sources. These significant efficiency gains indicate that $\kappa$-RRSS reduces computational overhead while maintaining reliable performance. Further comparisons of accuracy between w/ and w/o $\kappa$-RRSS across different number of sources and models can be found in Appendix E.3.
@@
# Src $\kappa$-RSS $\kappa$-RRSS $5$ $0.588$ $0.597$ $10$ $0.663$ $0.727$ $20$ $0.679$ $0.748$
Table 3: Ablation study on distortion of reliability estimation without $f_{\text{align}}$ on Llama3-8B-Instruct and TQA dataset.
Method Accuracy Oracle WMV $0.541$ Ours (w/) $0.537$ Ours (w/o) $0.490$
The importance of relevance in $\kappa$-RRSS. By comparing $\kappa$-RRSS (Reliable and Relevant Source Selection) and $\kappa$-RSS (Reliable Source Selection), we analyze the importance of incorporating relevance in source selection. Table 3 shows that incorporating relevance consistently improves accuracy, as high reliability alone does not ensure that sources contain documents relevant to the given query. This effect becomes more significant as the number of sources increases, providing a broader pool of relevant sources for selection.
Effectiveness of filtering with $f_{\textbf{align}}$. We evaluate the effectiveness of $f_{\text{align}}$ across three types of retrieved documents: factual, misinformation, and irrelevant. Table 4 shows the proportions of responses, both without (w/o) and with (w/) filtering, categorized by response types: correct, incorrect, IDK, and hallucination (i.e., not belonging to any other category). These results are based on 1,600 queries in the TQA dataset, using a single source with $p_{i}=0.5$ and $r_{i}=0.5$. Notably, without $f_{\text{align}}$, LLMs often generate correct (26.07%) or hallucinated responses (22.89%) that are not grounded in the retrieved documents when the retrieved documents are irrelevant. A similar trend is observed with misinformation documents. After applying $f_{\text{align}}$, these misaligned responses are substantially reduced (highlighted in blue), by replacing them with IDK (highlighted in red).
To illustrate this risk, we conduct experiments using the adversary-hammer prior, where four adversaries have $r_{i}=0.1$ and one hammer has $r_{i}=0.6$, utilizing Llama3-8B-Instruct and the TQA dataset. Due to the small $r_{i}$ of adversaries, which results in a lack of relevant documents, we use 800 queries for reliability estimation and the remaining 800 queries for test evaluation. As shown in Table 3, without filtering, the estimated weights become distorted, assigning more weight to adversaries and degrading performance. However, applying filtering effectively mitigates this distortion, bringing performance close to Oracle WMV.
Table 4: Answer type distribution (%) by retrieved document type in the filtering $f_{\text{align}}$ ablation study on Llama3-8B-Instruct model and TQA dataset. Additional results for other datasets and models are provided in Appendix E.5.
Figure 5: The results of reliability estimation under augmented variation for User A. Additional results for Politician A and B are provided in Appendix J.
@@ 6 Real-world Application
We demonstrate the practical applicability of our iterative reliability estimation method by collecting claims from real-world sources. The experimental setup is provided in Section 6.1, and the results are presented in Section 6.2.
@@ 6.1 Setup
Data collections. To collect real-world claims, we leverage a fact-checking platform PolitiFact that evaluates the truthfulness of claims requiring verification, such as those made by public figures. Specifically, we select two prominent public figures, Politician A and Politician B, as sources, collecting 388 claims (64 true, 324 false) and 104 claims (63 true, 41 false), respectively, using PolitiFact’s verdicts to determine their truthfulness. To further validate our method, we gather posts from social media, where unverified information spreads rapidly. We select User A, an account on X that shares breaking news, collecting 244 posts (180 true, 64 false) from January 1–13, 2025. We manually verify their truthfulness by cross-checking with fact-checking sites.
Experimental settings. We conduct experiments in two settings: (i) using the full set of collected real-world data, and (ii) augmenting the dataset by varying oracle reliability levels, adjusting the true-to-false ratio from 0.1 to 0.9.
Since the collected data from three sources may not fully capture diverse scenarios, we employ setting (ii) to evaluate our method under varying reliability conditions. Given the inherent challenge of fact-checking, this augmentation offers a scalable alternative for evaluating reliability estimation across different reliability levels.
Reliability estimation process. To apply our reliability estimation method, we generate yes/no fact-checking queries for each collected claim (e.g., “Is it true that {claim}?”), allowing for straightforward cross-checking of claims across multiple sources. We use Google News as a retriever to retrieve relevant documents, selecting the top-20 results, and GPT-4o-mini as the language model to generate responses.
@@ 6.2 Experimental results
7 Conclusion and Future Works
@@
We show that our reliability estimation method effectively estimates the reliability of real-world sources, particularly for news-related claims with abundant fact-checking sources. However, it remains challenging to apply to specialized topics (e.g., medical research) due to limited references for cross-verification. Exploring expert knowledge as an alternative could help address this limitation and presents a promising direction for future work. Additionally, since our framework operates in an offline setting, it requires periodic updates to adapt to the dynamic nature of source reliability. A promising direction for future work is the development of an online framework that continuously updates reliability estimates in real-time, enabling adaptive responses to the evolving information landscape.
Impact Statement
Our paper presents a framework that enhances the robustness of RAG in practical deployments by estimating source reliability. By improving the accuracy of information retrieval and generation, our approach helps mitigate factual errors in RAG systems. We hope these findings contribute to the development of more trustworthy RAG systems.
Appendix A Overview of RA-RAG
Figure 6: Overview of the RA-RAG framework. In the first stage, RA-RAG iteratively estimates the reliability of each source $v_{i}$ for $i\in[N]$ based on estimated true answers $\hat{y}^{j}$ for each question $j$, as outlined in equation (3) and (4). Based on the estimated reliability, in the second stage, the retriever selects $\kappa$ sources through the Reliable and Relevant Source Selection ($\kappa$-RRSS) process, detailed in Section 4.1. The final answer is determined using a weighted majority voting process, with the weights derived from the estimated reliability.
Appendix B Details of Misalignment Filtration
AlignScore (Zha et al., 2023) is a factual consistency evaluation method that assesses how well the generated text aligns with the given context. However, LLM outputs are often not in declarative sentences, and important contextual information is sometimes embedded in the query, making direct consistency evaluation challenging. To address this, we employ a sequence-to-sequence model from (Song, 2022), previously used in (Zha et al., 2023), to convert outputs into declarative sentences. Formally, we denote the declarative form of $\hat{y}_{i}$ as $\hat{y}_{i}^{*}$. With this conversion, the misalignment filtering process is as follows:
$f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))=\begin{cases}\text{IDK}&\text{if }\mathcal{E}(\hat{y}_{i};q,\mathcal{R}(q,\mathcal{S}_{i}))<\tau,\\ \hat{y}_{i}&\text{otherwise}.\end{cases}$ (5)
where $\mathcal{E}$ is the Alignscore function, $\mathcal{E}(\hat{y}_{i};q,\mathcal{R}(q,\mathcal{S}_{i}))=\mathcal{E}(\hat{y}_{i}^{*},\mathcal{R}(q,\mathcal{S}_{i}))$, and $\tau$ is the threshold. In all experiments, we set $\tau=0.1$ following Lei et al. (2023), which identifies this threshold as optimal for a real-world dataset comprising CNN and Daily Mail articles. For further analysis, we also conduct an ablation study on $\tau$ in Section E.5.
Appendix C $\kappa$-Reliable and Relevant Source Selection ($\kappa$-RRSS)

### v3

@@ 5 Experiments
Figure 2: Accuracy performance on NQ and TQA datasets. (a) Results with heterogeneous reliability via beta priors for varying sources (4 to 9) across the Llama3-8B-Instruct and GPT-4o-mini models. See Appendix F.1 on the HotpotQA dataset and Phi3-mini-Instruct model. (b) Results with adversarial setting via adversary-hammer prior for varying adversaries (1 to 7) with Llama3-8B-Instruct model, highlighting overall trends. Exact values, which may overlap significantly, are provided in Appendix A4 with HotpotQA results.
@@ 5.1 Experimental setups
Datasets. We construct a multi-source RAG benchmark with heterogeneous source reliability, using three question-answering (QA) datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQA) (Joshi et al., 2017), and HotpotQA (Yang et al., 2018). For each dataset, we generate both diverse factual documents and misinformation to simulate a source $S_{i}$ with varying reliability. Each source $\mathcal{S}_{i}$ is characterized by two parameters: reliability $p_{i}$, which represents the probability of providing factual information, and coverage $r_{i}$, which indicates the probability of containing relevant documents for a given query. To model source reliability $p_{i}$, we adopt two widely used priors from the reliability estimation literature Liu et al. (2012); Li and Yu (2014):
  • Beta prior: $p_{i}$ is sampled from $\text{Beta}\left(\nicefrac{{2\bar{w}}}{{1-\bar{w}}},\ 2\right)$ with an expected mean of $\bar{w}$. This setup reflects scenarios where sources exhibit a continuous spectrum of reliability, rather than strictly “reliable" or “unreliable". Following Liu et al. (2012); Li and Yu (2014), we set $\bar{w}=0.6$, balancing the presence of reliable and unreliable sources.
@@
For analytical simplicity, we set $r_{i}=0.6$ for both priors to focus on evaluating $p_{i}$. The details of the data generation and source construction processes are provided in Appendix H. Due to the computational and financial constraints, we use 1,600 queries per dataset, allocating 200 queries for reliability estimation and 1,400 queries for test evaluation.
Baselines. We compare our framework against RA-RAG and six baselines. (1) Oracle WMV assumes perfect knowledge of source reliability and directly uses these values as weights in Equation (2), representing the ideal scenario for multi-source RAG. (2) MV assigns equal weight to all sources, setting $v_{i}=1$ in Equation (2), disregarding source reliability. (3) Vanilla RAG Lewis et al. (2020) follows the standard RAG approach, retrieving documents without additional modules. (4) Robust RAG Xiang et al. (2024) is the first certifiably robust defense framework that enhances robustness by aggregating keywords from independent passages, assuming that the majority of retrieved documents are trustworthy. (5) Self-RAG Asai et al. (2024) is an advanced RAG that improves performance through adaptive retrieval, reducing irrelevant documents by leveraging specialized reflection tokens to improve factual accuracy. (6) Vanilla LLM generates responses without retrieval. Among these baselines, (1) and (2) are designed for multi-source RAG, while (3), (4), and (5) follow the standard RAG approach.
Models. For language models, we use Llama3-8B-Instruct (Dubey et al., 2024), Phi3-mini-Instruct (Abdin et al., 2024), GPT-4o-mini (OpenAI, 2024), and Llama2-7B (Touvron et al., 2023). As a retriever, we use Contriever (Izacard et al., 2022). Due to space limitations, the results for Llama2-7B with Self-RAG fine-tuned on Llama2-7B are provided in Appendix E.
@@
Figure 3: A qualitative example comparing the answers produces by MV and RA-RAG for a query from the NQ dataset. Additional examples are available in Appendix G.
Evaluation metric. Following prior works Mallen et al. (2023); Asai et al. (2024), we use accuracy as an evaluation metric, based on whether gold answers are included in model-generated responses. All results are averaged over $10$ random trials.
@@ 5.2 Main results
Beta prior. We evaluate RA-RAG across varying numbers of sources to assess its effectiveness in heterogeneous source reliability. As shown in Figure 2(a), RA-RAG consistently outperforms baselines, with performance gains increasing as more sources are incorporated. These results demonstrate the robustness of our approach in aggregating information from multiple sources with varying reliability. Notably, by selecting a subset of reliable and relevant sources using $\kappa$-RRSS, RA-RAG achieves performance comparable to Oracle WMV while improving efficiency by relying on fewer sources. In contrast, Robust RAG struggles with varying source reliability, as its certification assumption does not hold, resulting in lower performance than MV. Additionally, RA-RAG significantly outperforms Self-RAG, as shown in the Appendix E. These results emphasize the importance of differentiating between sources to prevent retrieval results from being overwhelmed by misinformation.
Figure 3 highlights the importance of considering source reliability when aggregating information across sources. While MV selects “November 2018" based only on response frequency, although it has low reliability, RA-RAG correctly identifies “November 2017" by leveraging well-estimated source reliabilities.
Adversary-hammer prior. To evaluate the robustness of RA-RAG in the worst-case scenario, we use the adversary-hammer prior with a total of 9 sources on the NQ dataset with Llama3-8B-Instruct, as shown in Figure 2(b). Our RA-RAG demonstrates significant robustness against adversaries, whereas Robust RAG and Vanilla RAG suffer severe performance degradation as the number of adversaries increases. Similarly, Self-RAG experiences significant performance degradation, as detailed in Appendix E. Notably, when the number of adversaries exceeds four, the performance of MV significantly degrades due to the dominance of misinformation, leading MV to select incorrect answers.
@@ 5.3 Ablation studies and analysis
Due to space limitations, we present the analysis of the effectiveness of filtering in Appendix C.
Impact of $\kappa$ for $\kappa$-RRSS. We conduct an ablation study on $\kappa$ across three datasets using Llama3-8B-Instruct with 9 sources. As shown in Figure 4, RA-RAG achieves stable performance starting from $\kappa=4$, indicating that selecting a small subset of reliable and relevant sources can maintain performance while significantly reducing computational overhead. This trend is consistent across other datasets; refer to Appendix F.2.
Computational efficiency of $\kappa$-RRSS. To assess the impact of $\kappa$-RRSS on computational efficiency, we compare RA-RAG in two configurations: with and without $\kappa$-RRSS. We measure four computational metrics: token consumption, API calls, inference cost, and wall-clock time. Specifically, token consumption refers to the total number of tokens processed per query during inference, including both input and output tokens. API calls measure the number of external API requests per query. Inference cost represents the computational expense (\$ per query) based on the GPT-4o-mini pricing policy.
As shown in Table 2, incorporating $\kappa$-RRSS consistently enhances computational efficiency across all metrics, with the reduction rate increasing as the number of sources grows. For example, in terms of token consumption, $\kappa$-RRSS reduces the total tokens processed by 2.6% with 5 sources, 32.3% with 10 sources, and 99.1% with 1000 sources. These significant efficiency gains indicate that $\kappa$-RRSS reduces computational overhead while maintaining reliable performance. Additional wall-clock time comparisons with baseline methods are provided in Appendix K. Further comparisons of accuracy between w/ and w/o $\kappa$-RRSS across different number of sources and models can be found in Appendix F.3.
The importance of relevance in $\kappa$-RRSS. To analyze the importance of incorporating relevance in $\kappa$-RRSS, we explore $\kappa$-Reliable Source Selection ($\kappa$-RSS), which chooses only the $\kappa$ most reliable sources without checking for relevance. Table 1 shows that incorporating relevance consistently improves accuracy, as high reliability alone does not ensure that sources contain documents relevant to the given query. This effect becomes more significant as the number of sources increases, providing a broader pool of relevant sources for selection.
# Src $\kappa$-RSS $\kappa$-RRSS $5$ $0.588$ $0.597$ $10$ $0.663$ $0.727$ $1000$ $0.689$ $0.768$
@@ 6 Real-world Application
We demonstrate the practical applicability of our iterative reliability estimation method by applying it to real-world sources. The experimental setup is described in Section 6.1, and the results are presented in Section 6.2.
# Src $\kappa$-RRSS Token Consumption ($\downarrow$) API Calls ($\downarrow$) Inference Cost ($\downarrow$) Accuracy ($\uparrow$) $5$ w/o $3138$ $5$ $0.00048$ $0.597$ w/ $3055$ ($\downarrow 2.6\%$) $4.87$ ($\downarrow 2.6\%$) $0.00046$ ($\downarrow 4.2\%$) $0.597$ $10$ w/o $6272$ $10$ $0.00096$ $0.744$ w/ $4251$ ($\downarrow 32.3\%$) $6.79$ ($\downarrow 32.1\%$) $0.00065$ ($\downarrow 32.3\%$) $0.727$ $1000$ w/o $627115$ $1000$ $0.096$ $0.780$ w/ $5415$ ($\downarrow 99.1\%$) $8.66$ ($\downarrow 99.1\%$) $0.00083$ ($\downarrow 99.1\%$) $0.768$
Table 2: Comparison of computational efficiency with and without $\kappa$-RRSS evaluated on the NQ dataset using GPT-4o-mini. The reported values represent the average per query, with the values in parentheses ()($\cdot$)( ⋅ ) indicating the reduction rate achieved with $\kappa$-RRSS.
Figure 4: Accuracy across different values of $\kappa$ on Llama3-8B-Instruct and NQ dataset. Results for other datasets are provided in Appendix F.2
@@ 6.1 Setup
Data collections. To collect real-world claims, we leverage a fact-checking platform PolitiFact that evaluates the truthfulness of claims requiring verification, such as those made by public figures. Specifically, we select two prominent public figures, Politician A and Politician B, as sources, and collect 388 claims (64 true, 324 false) and 104 claims (63 true, 41 false), respectively, using PolitiFact’s verdicts to determine their truthfulness. As an alternative information source, we gather posts from social media, where unverified information spreads rapidly. We select User A, an account on X that shares breaking news, collecting 244 posts (180 true, 64 false) from January 1-13, 2025. We manually verify their truthfulness by cross-checking with fact-checking sites.
Experimental settings. We conduct experiments in two settings: (i) using the full set of collected real-world data, and (ii) augmenting the dataset by varying oracle reliability levels, adjusting the true-to-false ratio from 0.1 to 0.9 through random sampling.
As the data collected from the three sources may not fully capture the diversity of real-world scenarios, we include setting (ii) to evaluate our method under a broader range of reliability conditions. Given the inherent challenge of fact-checking, this augmentation offers a scalable alternative for evaluating reliability estimation across different reliability levels.
Reliability estimation process. To apply our reliability estimation method, we generate yes/no fact-checking queries for each collected claim (e.g., “Is it true that {claim}?"), allowing for straightforward cross-checking of claims across multiple sources. We use Google News as the retriever, which provides rich sources for retrieving relevant documents by selecting the top 20 results. GPT-4o-mini is used as the language model to generate responses.
@@ 6.2 Experimental results
Figure 5: Results of reliability estimation under augmented variation for User A. Additional results for Politician A and B are in Appendix L.
7 Conclusion
@@
8 Limitations
We show that our reliability estimation method effectively estimates the reliability of real-world sources, particularly for news-related claims with abundant fact-checking sources. However, it remains challenging to apply to specialized topics due to limited references for cross-verification. Exploring expert knowledge as an alternative could help address this limitation and presents a promising direction for future work. Additionally, since our framework operates in an offline setting, it requires periodic updates to capture changes in source reliability over time. While such updates can be performed efficiently, as demonstrated in Appendix K.2, there remains a risk that even reliable sources may suddenly disseminate large volumes of unreliable information, such as in cases of account hacking. Although such cases are uncommon, this threat underscores the need for more responsive systems. A promising direction for future work is the development of an online framework that continuously updates reliability estimates in real time, enabling adaptive responses to the evolving information landscape.
Appendix A Details of Misalignment Filtration
AlignScore Zha et al. (2023) is a factual consistency evaluation method that assesses how well the generated text aligns with the given context. However, LLM outputs are often not in declarative sentences, and important contextual information is sometimes embedded in the query, making direct consistency evaluation challenging. To address this, we employ a sequence-to-sequence model from Song (2022), previously used in Zha et al. (2023), to convert outputs into declarative sentences. Formally, we denote the declarative form of $\hat{y}_{i}$ as $\hat{y}_{i}^{*}$. With this conversion, the misalignment filtering process is as follows:
$f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))$
=$=\begin{cases}\text{IDK}&\text{if }\mathcal{E}(\hat{y}_{i};q,\mathcal{R}(q,\mathcal{S}_{i}))<\tau\\ \hat{y}_{i}&\text{otherwise}\end{cases}\;,$
where $\mathcal{E}$ is the Alignscore function, $\mathcal{E}(\hat{y}_{i};q,\mathcal{R}(q,\mathcal{S}_{i}))=\mathcal{E}(\hat{y}_{i}^{*},\mathcal{R}(q,\mathcal{S}_{i}))$, and $\tau$ is the threshold. In all experiments, we set $\tau=0.1$ following Lei et al. (2023), which identifies this threshold as optimal for a real-world dataset comprising CNN and Daily Mail articles. For further analysis, we also conduct an ablation study on $\tau$ in Section M.
Appendix B $\kappa$-Reliable and Relevant Source Selection ($\kappa$-RRSS)